OpenStudy (anonymous):

how to find x and y from equation 2x +3y =-3 3x-2y=28? please help :D

OpenStudy (hoblos):

2x +3y =-3 *2 3x-2y=28 *3 4x +6y= -6 9x -6y =84 add them 13x= 78 x=6 => 2(6) +3y = -3 => 3y= -15 y= -5

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

my pleasure XD